1) It is known that in a certain manufacturing process, 1% of the products are defective. If a customer purchases 50 of these products selected at random what is probability that he receives 2 or less defective products? Epts.2) F items selected at random what is the expected value and the standard deviation of the number of defects? or the manufacturing process in Problem 1, if another customer purchases 200
Answer #1

Solution:

probability of finding defective product p = 0.01

probability of not finding a defective product q = 1 - 0.01 = 0.99

1) number of products customer purchased n = 50

Probability that he receives 2 or less defective products P(X \leq 2) = P(X = 2) + P(X = 1) + P(X = 0)

Formula : P(X = k) = n!/ [k! (n-k)!] (p)k (q)n - k

  P(X \leq 2) = 0.0756 +0.3056 +0.6050

= 0.9862

2) number of products customer purchased n = 200

expected value = 0* P(X =0) + 1* P(X = 1) + 2 * P(X = 2) + ...................+ 200 * P(X = 200)

= 0* (0.1340) + 1 * (0.2707) + 2 * (0.2720) + ..................+ 200 *(0)

= 1.9997

Standard deviation \sigma= \sqrt{\sum x^2 . p(x) - \mu ^2}

\sumx2 . p(x) = 02 * (0.1340) + 12 * (0.2707) +...............+ 2002 * (0)

= 5.9773

\mu^2 = 0.98622 = 0.9725

  \sigma= \sqrt{\sum x^2 . p(x) - \mu ^2} = \sqrt{5.9773 - 0.9725}

  \approx1.4066